OpenStudy (anonymous):

Write 4.48 as the quotient of two integers. A. 75./336 B. 118./25 C. 336/75 D. 336/150

OpenStudy (anonymous):

@Psymon

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

divide each one of the answers until you get 4.48

OpenStudy (anonymous):

C

OpenStudy (anonymous):

4.48=448/100=336/75